OpenStudy (anonymous):

What is the angular momentum of the disk as it leaves the plane and rolls on a level surface? The radius of the disk is 0.1 m.

OpenStudy (anonymous):

it rolls down an incline? how far does it fall?

OpenStudy (anonymous):

I agree with Stan, it's not a clear statement.

OpenStudy (anonymous):

Well there is a previous problem that ties in with this problem. I just realized it. My instructor didn't make this very clear to us. The problem..."A 500 g solid disk rolls down an inclined plane 1 m long elevated by 10 cm. What is the kinetic energy of the disk at the bottom of the plane?" "Solution: 1/2(9.8)(0.1) = 0.49 J". And then the question I asked ties in with this problem I just provided.

OpenStudy (anonymous):

I assume he wants the angular momentum about the center of the disk...\[L=I \omega\]where the angular velocity wanted in the angular velocity at the bottom of the hill. You can get that from \[K = K _{rotation} + K _{translation}\]\[0.49J=1/2I \omega^{2}+1/2mv^{2}\]\[I=1/2mR^{2}\]and assuming rolling w/o slipping \[v=r \omega\]so\[0.49J=1/2(1/2mR^{2})\omega^{2}+1/2m(R^{2}\omega^{2})\]solve for omega and substitute into the equation for ang. momentum

OpenStudy (anonymous):

the change in potential energy is already taken care of in the gain of kinetic energy (0.49J).... you don't want to do it twice :)

OpenStudy (anonymous):

Sorry ! Exact ! I would be tired. So, I will supress my comment.

OpenStudy (anonymous):

Just for my curiosity, is really the question : "What is the kinetic energy of the disk at the bottom of the plane?" => Kinetic energy or potential energy. I just try to find a mitigating circumstance. :)

OpenStudy (anonymous):

The previous question (the lead-in to this one) wanted to find the total KE, which you get from the potential energy, but here they want L